OpenStudy (anonymous):

2x – y + z = –1 x – y + z = 1 x – 2y + z = 2 x=? y=? z=?

OpenStudy (cwrw238):

2x – y + z = –1 x – y + z = 1 subtract x = -2 so we have -2 - y + z = 1 -2 -2y + z = 2 subtract y = -1 from first equation -4 -(-1) + z = -1 can you finish this off

OpenStudy (anonymous):

Yes, I got it, thank you